OpenStudy (anonymous):

Write as a sum of partial fractions. x/(3x-2)(2x+1)

hartnn (hartnn):

\(\large \dfrac{1}{(ax+b)(cx+d)}=\dfrac{A}{ax+b}+\dfrac{B}{cx+d}\)

OpenStudy (anonymous):

so is it A/(2x+1) + B/(3x-2) ?

hartnn (hartnn):

thats absolutely correct :)

hartnn (hartnn):

do you also need to find values of A and B ?

OpenStudy (shiraz14):

Yes - substitute x=2/3 and x=-0.5 in the original equation to find A & B.

OpenStudy (shiraz14):

You'll have to do this one at a time.

OpenStudy (anonymous):

when you do that is one A and one B?

hartnn (hartnn):

yes, to get the value of A, you put x =-1/2 to get the value of B, you put x= 2/3